# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v25

@@ 3.17 Phonological Tools
On the case of phonological layer, a report on Sinhala phonetics and phonology was published by Wasala and Gamage [351]Wickramasinghe et al. [352] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems. The Massively Multilingual Speech (MMS)138138138https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al. [145] has Sinhala data for the Spoken Language Identification (LID) task. The meta-study by Al-Fraihat et al. [353] compared the status of Sinhala speech recognition research against 17 other languages.
@@ 3.17.1 Text-to-Speech
Based on the earlier work by Weerasinghe et al. [354]Wasala et al. [355] have developed methods for Sinhala grapheme-to-phoneme conversion along with a set of rules for schwa epenthesis. This work was then extended by Nadungodage et al. [356]Weerasinghe et al. [357] developed a Sinhala text-to-speech system. However, it is not publicly accessible. They internally extended it to create a system capable of helping a mute person achieve synthesized real-time interactive voice communication in Sinhala [358]. A rule based approach for automatic segmentation of a small set of Sinhala text into syllables was proposed by Kumara et al. [359]. An ew prosodic phrasing method to help with Sinhala Text-to-Speech process was proposed by Bandara et al. [360, 361, 362]Sodimana et al. [363] proposed a text normalization methodology for Sinhala text-to-speech systems. Further, Sodimana et al. [364] formalized a step-by-step process for building text-to-speech voices for Sinhala. Both Jayamanna [365] and Mishangi [366] have created Sinhala document readers for visually impaired persons to be used on Android devices. An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [367]. The works by Lakmal et al. [368] and Senarathna et al. [369] adapted MaryTTS [370] to synthesize Sinhala speech. The study by Jayawardhana et al. [371] used Deep Voice [372] for Sinhala and English TTS. Gamage et al. [373] included a Sinhala text-to-speech module as one of the three modules present in their currency recognition system. The study by Madhusha et al. [374] claims to have created a mobile app with Sinhala Text-to-Speech and OCR to read books for visually impaired students. Praveen et al. [375] claim to have created a Sinhala TTS system with a custom neural network that archives 98% accuracy.  Anuradha and Thelijjagoda [376] proposed a machine translation system to convert Sinhala and English Braille documents into voice. A separate group has done work on Sinhala text-to-speech systems independent to above [377].
@@ 3.17.2 Speech-to-Text
On the converse, Nadungodage et al. [378] have done a series of work on Sinhala speech recognition with special notice given to Sinhala being a resource poor language. This project divides its focus on: continuity [379], active learning [380], and speaker adaptation [381]. A Sinhala speech recognition for voice dialing which is speaker independent was proposed by Amarasingha and Gamini [382] and on the other end, a Sinhala speech recognition methodology for interactive voice response systems, which are accessed through mobile phones was proposed by Manamperi et al. [383]. A Sinhala speech to Unicode text converter for the disaster relief domain was proposed by Prasangini and Nagahamulla [384]Priyadarshani [385] proposes a method for speaker dependant speech recognition based on their previous work on: dynamic time warping for recognizing isolated Sinhala words [386], genetic algorithms [387], and syllable segmentation method utilizing acoustic envelopes [388]. The method proposed by Gunasekara and Meegama [389] utilizes an HMM model for Sinhala speech-to-text. A Sinhala speech recognizer supporting bi-directional conversion between Unicode Sinhala and phonetic English was proposed by Punchimudiyanse and Meegama [390]. The work by Karunanayake et al. [391] transfer learns CNNs for transcribing free-form Sinhala and Tamil speech data sets for the purpose of classification. Dilshan [392] conducted a study for the specific use case of transcribing number sequences in continuous Sinhala speech.  Gamage et al. [393] explored the use of combinational acoustic models such as Deep Neural Network - Hidden Markov Model (DNN-HMM) [394] and Subspace Gaussian Mixture Model (SGMM) [394] in Sinhala speech recognition. In the follow-up work, Gamage et al. [395] extended that work with end-to-end Lattice-Free Maximum Mutual Information (e2e LF-MMI) model [396] which is claimed to be a viable solution for low resource language speech recognition by Carmantini et al. [397]. However, it was shown that the new model slightly underperforms compared to the state-of-the-art result. Later, they conducted further development on their model in a follow-up work [398].  Karunathilaka et al. [399] explore Sinhala speech recognition using deep learning models such as: pre-trained DNN, DNN, TDNN, TDNN+LSTM. The first half of the study by Arafath [400] dealt with recognizing Sinhala speech using LSTMs. Gamage et al. [373] included a Sinhala speech recognition module as one of the three modules present in their currency recognition system. Time-delay neural network architectures (including multistream CNN architecture) were used for acoustic modeling of Sinhala Automatic Speech Recognition (ASR) by Warusawithana et al. [401]. They have used the Kaldi speech recognition toolkit [402] for training the ASR models. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [403], proposed a Sinhala speech recognition system. The study by Azir et al. [404] attempts to identify number sequences spoken in Sinhala. TacoSi introduced by Arachchige and Weerasinghe [405] is based on Tacotron [406] and has been evaluated with 10 human evaluators to determine its text-to-speech quality. Nanayakkara [407], Nanayakkara and Weerasinghe [408] used DeepSpeech139139139https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition. Gunarathne et al. [409] used an earlier version of the CMUSphinx toolkit140140140https://cmusphinx.github.io/ to transcribe Sinhala speech to text. The later work by Akesh and Meegama [410] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [411, 412] to automatically generate Sinhala subtitles from Speech. The work by Wickramaarachchi et al. [413] also used MFCC features but focused on Sinhala speech intonation recognition for the purpose of detecting speech impediments in young children. Dissanayaka et al. [414] have introduced Word Sri an application that is voice-activated and capable of Sinhala grammar checking and plagiarism checking. tha [415] have introduced a Sinhala text to Speech dataset141141141https://github.com/SiTa-SpeakerDiarization/SiTa transcribed from 60 videos with a total of 602 minutes. As the baseline, they report the lowest Diarization Error Rate (DER) of 6.0 for Powerset Cross-Entropy Diarization [416]. While the data set link points to a github repository, it does not contain the data set; it points to a project website142142142https://sita-speakerdiarization.github.io/ which then points to a download page143143143https://rtuthaya.staff.uom.lk/resources/dataset/44 where you need to send a request to the authors to obtain the data set.
@@ 3.17.3 Speech-to-Speech
Layansan et al. [417] created a speech-to-speech translation system for Sinhala on the Android platform. The system developed by Rajapakshe et al. [418] is also speech-to-speech in the sense that, it is a chatbot for scheduling medical appointments and giving medical advice where the front end contains speech recognition and voice synthesizer components that interfaced with a chatbot component in the back end. The work by Athas and Pirapuraj [419] uses Google APIs (speech-to-text API, text-to-text-translator API, and text-to-speech synthesizer API) to achieve voice translation for Sinhala and Tamil.
@@ 3.17.5 Speech classification
Luger et al. [423] introduced a data set named unsupervised_peoples_speech144144144https://huggingface.co/datasets/MLCommons/unsupervised_peoples_speech that conatains 981 hours of Sinhala speech data. The Sinhala speech classification system proposed by Buddhika et al. [248] does so without converting the speech-to-text. However, they report that this approach only works for specific domains with well-defined limited vocabularies. The work by Dinushika et al. [424] uses automatic speech recognition of Sinhala for speech command classification. Extending that, Kavmini et al. [425] presented a Sinhala speech command classification system which can be used for downstream applications. The voice assistant system created by Senarathne et al. [426] is capable of handling Sinhala voice commands. Kathriarachchi and Pulasinghe [427] developed a system to classify and predict Sinhala speech clips of children into age groups. The work by Welarathna et al. [428] used CNNs to classify emotions (sad, disgust, surprise, neutral, happy, calm, fear, and angry) of Sinhala speech by Autistic children. Sundarapperuma [429] created a speech emotion detection system for Sinhala.
@@ 3.18 Optical Character Recognition Applications
While it is not necessarily a component of the NLP stack shown in Fig 3, which follows the definition by Liddy [36], it is possible to swap out the bottom most phonological layer of the stack in favour of an Optical Character Recognition (OCR) and text rendering layer.
@@ 3.18.1 Printed Text
The XTREME-UP145145145https://github.com/google-research/xtreme-up data set created by Ruder et al. [144] contains a Sinhala data set for the OCR task. The data was obtained from book transcriptions. An attempt for Sinhala OCR system has been taken by Rajapakse et al. [433] before any other work has been done on the topic. Much later, a linear symmetry-based approach was proposed by Premaratne and Bigun [434, 435]. They then used hidden Markov model-based optimization on the recognized Sinhala script [436]. Similarly, Hewavitharana et al. [437] used hidden Markov models for off-line Sinhala character recognition. Herath et al. [123], Herath and Medagoda [124] developed a prepossessing engine based on a template matching for printed Sinhala OCR. Statistical approaches with histogram projections for Sinhala character recognition is proposed by Hewavitharana and Kodikara [438], by Ajward et al. [439], and by Madushanka et al. [440].  Karunanayaka et al. [441] also did off-line Sinhala character recognition with a use case for postal city name recognition. A separate group had attempted Sinhala OCR [442] mainly involving the nearest-neighbor method [443, 444]. A study by Ediriweera [445] uses dictionaries to correct errors in Sinhala OCR. An early attempt for Sinhala OCR by Dias et al. [446] has been extended to be online and made available to use via desktops [447] and hand-held devices [448] with the ability to recognize handwriting. A simple neural network based approach for Sinhala OCR was utilized by Rimas et al. [449]. A fuzzy-based model for identifying printed Sinhala characters was proposed by Gunarathna et al. [450]Premachandra et al. [451] proposes a simple back-propagation artificial neural network with hand crafted features for Sinhala character recognition. Another neural network with specialized feature extraction for Sinhala character recognition was proposed by Jayamaha and Naleer [452]. On the matter of neural networks and feature extraction, a feature selection process for Sinhala OCR was proposed by Kumara and Ragel [453].  Jayawickrama et al. [454] worked on Sinhala printed characters with special focus on handling diacritic vowels. However, they opted to refer to diacritic vowels as modifiers in their work. Gunawardhana and Ranathunga [455] proposed a limited approach to recognize Sinhala letters on Facebook images. A CNN-based methodology to improve printed Sinhala character OCR was proposed by Liyanage [456]. Printed Character Recognition (PCR) was used by Vasantharajan and Thayasivam [113] to create a large-scale Tamil-Sinhala-English parallel corpus. A meta-study on the effects of text genre, image resolution, and algorithmic complexity needed for Sinhala OCR from books and newspapers was conducted by Anuradha et al. [457].  Anuradha et al. [458] used Tesseract 3146146146https://tesseract-ocr.github.io/ [459] for Sinhala OCR. The later study by [460] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.  Maduranga and Jayalal [461] used Artificial Neural Network (ANN) based on Universe of Discourse and Self Organization Map methods to recognize multi-style printed Sinhala characters. The study by De Silva et al. [462], even though ostensibly presented as translation research, is just invoking Google Cloud service for translation. Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents. The work by Thamarasee and Surendra [463] used histogram-oriented gradient descriptor (HOG descriptor) [464] and SVMs to recognize 400+ variations of Sinhala characters.

### v26

@@ 3.17 Phonological Tools
On the case of phonological layer, a report on Sinhala phonetics and phonology was published by Wasala and Gamage [369]Wickramasinghe et al. [370] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems. The Massively Multilingual Speech (MMS)163163163https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al. [148] has Sinhala data for the Spoken Language Identification (LID) task. The meta-study by Al-Fraihat et al. [371] compared the status of Sinhala speech recognition research against 17 other languages.
@@ 3.17.1 Text-to-Speech
Based on the earlier work by Weerasinghe et al. [372]Wasala et al. [373] have developed methods for Sinhala grapheme-to-phoneme conversion along with a set of rules for schwa epenthesis. This work was then extended by Nadungodage et al. [374]Weerasinghe et al. [375] developed a Sinhala text-to-speech system. However, it is not publicly accessible. They internally extended it to create a system capable of helping a mute person achieve synthesized real-time interactive voice communication in Sinhala [376]. A rule based approach for automatic segmentation of a small set of Sinhala text into syllables was proposed by Kumara et al. [377]. An ew prosodic phrasing method to help with Sinhala Text-to-Speech process was proposed by Bandara et al. [378, 379, 380]Sodimana et al. [381] proposed a text normalization methodology for Sinhala text-to-speech systems. Further, Sodimana et al. [382] formalized a step-by-step process for building text-to-speech voices for Sinhala. Both Jayamanna [383] and Mishangi [384] have created Sinhala document readers for visually impaired persons to be used on Android devices. An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [385]. The works by Lakmal et al. [386] and Senarathna et al. [387] adapted MaryTTS [388] to synthesize Sinhala speech. The study by Jayawardhana et al. [389] used Deep Voice [390] for Sinhala and English TTS. Gamage et al. [391] included a Sinhala text-to-speech module as one of the three modules present in their currency recognition system. The study by Madhusha et al. [392] claims to have created a mobile app with Sinhala Text-to-Speech and OCR to read books for visually impaired students. Praveen et al. [393] claim to have created a Sinhala TTS system with a custom neural network that archives 98% accuracy. The work by Senarath [394] discusses using existing TTS systems for Sinhala. They especially highlight VAENAR-TTS [395] as the best candidate for this task. Nayanathara et al. [396], on the other hand, propose to use the VITS architecture [397]164164164https://github.com/jaywalnut310/vits for the same task.  Anuradha and Thelijjagoda [398] proposed a machine translation system to convert Sinhala and English Braille documents into voice. A separate group has done work on Sinhala text-to-speech systems independent to above [399].
@@ 3.17.2 Speech-to-Text
On the converse, Nadungodage et al. [400] have done a series of work on Sinhala speech recognition with special notice given to Sinhala being a resource poor language. This project divides its focus on: continuity [401], active learning [402], and speaker adaptation [403]. A Sinhala speech recognition for voice dialing which is speaker independent was proposed by Amarasingha and Gamini [404] and on the other end, a Sinhala speech recognition methodology for interactive voice response systems, which are accessed through mobile phones was proposed by Manamperi et al. [405]. A Sinhala speech to Unicode text converter for the disaster relief domain was proposed by Prasangini and Nagahamulla [406]Priyadarshani [407] proposes a method for speaker dependant speech recognition based on their previous work on: dynamic time warping for recognizing isolated Sinhala words [408], genetic algorithms [409], and syllable segmentation method utilizing acoustic envelopes [410]. The method proposed by Gunasekara and Meegama [411] utilizes an HMM model for Sinhala speech-to-text. A Sinhala speech recognizer supporting bi-directional conversion between Unicode Sinhala and phonetic English was proposed by Punchimudiyanse and Meegama [412]. The work by Karunanayake et al. [413] transfer learns CNNs for transcribing free-form Sinhala and Tamil speech data sets for the purpose of classification. Dilshan [414] conducted a study for the specific use case of transcribing number sequences in continuous Sinhala speech.  Gamage et al. [415] explored the use of combinational acoustic models such as Deep Neural Network - Hidden Markov Model (DNN-HMM) [416] and Subspace Gaussian Mixture Model (SGMM) [416] in Sinhala speech recognition. In the follow-up work, Gamage et al. [417] extended that work with end-to-end Lattice-Free Maximum Mutual Information (e2e LF-MMI) model [418] which is claimed to be a viable solution for low resource language speech recognition by Carmantini et al. [419]. However, it was shown that the new model slightly underperforms compared to the state-of-the-art result. Later, they conducted further development on their model in a follow-up work [420].  Karunathilaka et al. [421] explore Sinhala speech recognition using deep learning models such as: pre-trained DNN, DNN, TDNN, TDNN+LSTM. The first half of the study by Arafath [422] dealt with recognizing Sinhala speech using LSTMs. Gamage et al. [391] included a Sinhala speech recognition module as one of the three modules present in their currency recognition system. Time-delay neural network architectures (including multistream CNN architecture) were used for acoustic modeling of Sinhala Automatic Speech Recognition (ASR) by Warusawithana et al. [423]. They have used the Kaldi speech recognition toolkit [424] for training the ASR models. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [425], proposed a Sinhala speech recognition system. The study by Azir et al. [426] attempts to identify number sequences spoken in Sinhala. TacoSi introduced by Arachchige and Weerasinghe [427] is based on Tacotron [428] and has been evaluated with 10 human evaluators to determine its text-to-speech quality. Nanayakkara [429], Nanayakkara and Weerasinghe [430] used DeepSpeech165165165https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition. Gunarathne et al. [431] used an earlier version of the CMUSphinx toolkit166166166https://cmusphinx.github.io/ to transcribe Sinhala speech to text. The later work by Akesh and Meegama [432] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [433, 434] to automatically generate Sinhala subtitles from Speech. The work by Wickramaarachchi et al. [435] also used MFCC features but focused on Sinhala speech intonation recognition for the purpose of detecting speech impediments in young children. Dissanayaka et al. [436] have introduced Word Sri an application that is voice-activated and capable of Sinhala grammar checking and plagiarism checking. tha [437] have introduced a Sinhala text to Speech dataset167167167https://github.com/SiTa-SpeakerDiarization/SiTa transcribed from 60 videos with a total of 602 minutes. As the baseline, they report the lowest Diarization Error Rate (DER) of 6.0 for Powerset Cross-Entropy Diarization [438]. While the data set link points to a github repository, it does not contain the data set; it points to a project website168168168https://sita-speakerdiarization.github.io/ which then points to a download page169169169https://rtuthaya.staff.uom.lk/resources/dataset/44 where you need to send a request to the authors to obtain the data set. Omnilingual ASR Team et al. [439] includes Sinhala170170170https://huggingface.co/datasets/facebook/omnilingual-asr-corpus/tree/main/data/sin_Sinh among the 1600+ languages for which they have created a multilingual ASR system171171171https://github.com/facebookresearch/omnilingual-asr.
@@ 3.17.3 Speech-to-Speech
Layansan et al. [440] created a speech-to-speech translation system for Sinhala on the Android platform. The system developed by Rajapakshe et al. [441] is also speech-to-speech in the sense that, it is a chatbot for scheduling medical appointments and giving medical advice where the front end contains speech recognition and voice synthesizer components that interfaced with a chatbot component in the back end. The work by Athas and Pirapuraj [442] uses Google APIs (speech-to-text API, text-to-text-translator API, and text-to-speech synthesizer API) to achieve voice translation for Sinhala and Tamil. Dilshani and Chandrasena [443] have conducted a survey on Speech-to-Speech systems available for Sinhala and Tamil translation. In their speech-to-speech chatbot, Weerakoon et al. [444] uses Wav2Vec2-BERT [445] for ASR and VITS [397]172172172https://github.com/jaywalnut310/vits model for TTS. The Text-to-Text chatbot component is the middle is handled by an LLM with RAG.
@@ 3.17.5 Speech classification
Luger et al. [449] introduced a data set named unsupervised_peoples_speech173173173https://huggingface.co/datasets/MLCommons/unsupervised_peoples_speech that conatains 981 hours of Sinhala speech data. The Sinhala speech classification system proposed by Buddhika et al. [259] does so without converting the speech-to-text. However, they report that this approach only works for specific domains with well-defined limited vocabularies. The work by Dinushika et al. [450] uses automatic speech recognition of Sinhala for speech command classification. Extending that, Kavmini et al. [451] presented a Sinhala speech command classification system which can be used for downstream applications. The voice assistant system created by Senarathne et al. [452] is capable of handling Sinhala voice commands. Kathriarachchi and Pulasinghe [453] developed a system to classify and predict Sinhala speech clips of children into age groups. The work by Welarathna et al. [454] used CNNs to classify emotions (sad, disgust, surprise, neutral, happy, calm, fear, and angry) of Sinhala speech by Autistic children. Sundarapperuma [455] created a speech emotion detection system for Sinhala.
@@ 3.18 Optical Character Recognition Applications
While it is not necessarily a component of the NLP stack shown in Fig 3, which follows the definition by Liddy [38], it is possible to swap out the bottom-most phonological layer of the stack in favour of an Optical Character Recognition (OCR) and text rendering layer. Gunathilaka et al. [459] introduced SinOCR174174174https://github.com/SriDoc/datasets, consisting of 100,000 images which include 1,135 handwritten Sinhala texts175175175https://drive.google.com/file/d/16X6lAzinZTIgiy6iw8Z0mLEyPyEwZJkL/view?usp=drive_link and texts printed in 200 different Sinhala fonts176176176https://drive.google.com/file/d/1uuPNN8M4jHLuZLF94CApkiG3kysDALia/view?usp=drive_link. They also released SinFUND177177177https://drive.google.com/file/d/1hfAfCJcZciEqMgO_wCuvTQ4pWI8t0L5D/view?usp=drive_link, a fully annotated dataset of 100 manually filled Sinhala forms.
@@ 3.18.1 Printed Text
The XTREME-UP178178178https://github.com/google-research/xtreme-up data set created by Ruder et al. [147] contains a Sinhala data set for the OCR task. The data was obtained from book transcriptions. An attempt for Sinhala OCR system has been taken by Rajapakse et al. [460] before any other work has been done on the topic. Much later, a linear symmetry-based approach was proposed by Premaratne and Bigun [461, 462]. They then used hidden Markov model-based optimization on the recognized Sinhala script [463]. Similarly, Hewavitharana et al. [464] used hidden Markov models for off-line Sinhala character recognition. Herath et al. [126], Herath and Medagoda [127] developed a prepossessing engine based on a template matching for printed Sinhala OCR. Statistical approaches with histogram projections for Sinhala character recognition is proposed by Hewavitharana and Kodikara [465], by Ajward et al. [466], and by Madushanka et al. [467].  Karunanayaka et al. [468] also did off-line Sinhala character recognition with a use case for postal city name recognition. A separate group had attempted Sinhala OCR [469] mainly involving the nearest-neighbor method [470, 471]. A study by Ediriweera [472] uses dictionaries to correct errors in Sinhala OCR. An early attempt for Sinhala OCR by Dias et al. [473] has been extended to be online and made available to use via desktops [474] and hand-held devices [475] with the ability to recognize handwriting. A simple neural network based approach for Sinhala OCR was utilized by Rimas et al. [476]. A fuzzy-based model for identifying printed Sinhala characters was proposed by Gunarathna et al. [477]Premachandra et al. [478] proposes a simple back-propagation artificial neural network with hand crafted features for Sinhala character recognition. Another neural network with specialized feature extraction for Sinhala character recognition was proposed by Jayamaha and Naleer [479]. On the matter of neural networks and feature extraction, a feature selection process for Sinhala OCR was proposed by Kumara and Ragel [480].  Jayawickrama et al. [481] worked on Sinhala printed characters with special focus on handling diacritic vowels. However, they opted to refer to diacritic vowels as modifiers in their work. Gunawardhana and Ranathunga [482] proposed a limited approach to recognize Sinhala letters on Facebook images. A CNN-based methodology to improve printed Sinhala character OCR was proposed by Liyanage [483]. Printed Character Recognition (PCR) was used by Vasantharajan and Thayasivam [116] to create a large-scale Tamil-Sinhala-English parallel corpus. A meta-study on the effects of text genre, image resolution, and algorithmic complexity needed for Sinhala OCR from books and newspapers was conducted by Anuradha et al. [484].  Anuradha et al. [485] used Tesseract 3179179179https://tesseract-ocr.github.io/ [486] for Sinhala OCR. The later study by [487] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.  Maduranga and Jayalal [488] used an Artificial Neural Network (ANN) based on Universe of Discourse and Self-Organisation Map methods to recognise multi-style printed Sinhala characters. The study by De Silva et al. [489], even though ostensibly presented as translation research, is just invoking Google Cloud service for translation. Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents. The work by Thamarasee and Surendra [490] used a histogram-oriented gradient descriptor (HOG descriptor) [491] and SVMs to recognise 400+ variations of Sinhala characters. After conducting a survey of existing methods [492], Hulathdoowage and Kumara [493] proposed a U-Net [494] based system for Sinhala printed document layout analysis. For this task they have created a data set using Sinhala textbooks. However, they do not provide a public link to this data set. Jayatilleke and de Silva [495] conducted a comparative analysis on zero-shot Sinhala OCR using a synthetic Sinhala OCR data set180180180https://huggingface.co/datasets/Ransaka/sinhala_synthetic_ocr-large. They concluded that Surya181181181https://github.com/VikParuchuri/surya to be the best Sinhala OCR system in this configuration. This aligns with the previous short study conducted by Velayuthan and Ambegoda [496]. However, later in their subsequent work, Jayatilleke and de Silva [497], they showed that for real-world applications Google Document AI182182182https://cloud.google.com/document-ai/ obtains better results for Sinhala.